Encino Plaza, Inc. owns a 50-acre shopping strip in Encino, California. In January 1994, an earthquake hit the area and destroyed parts of the shopping center. Some relevant information include the following:

Asset

Basis

FMV Before

FMV After

Insurance Recovery

Building

$90,000

$70,000

$0

$70,000

Equipment

$40,000

$50,000

$0

$25,000

Parking Structure

$90,000

$120,000

$70,000

$25,000

Because of the extensive damage caused by the earthquake, the President of the U.S. designated the area as a disaster area. Encino Plaza Inc. had $90,000 of taxable income last year. The company’s taxable income for the current year, excluding the loss from the earthquake, is $220,000. Determine the amount of the corporation’s loss and the year in which it should take the loss.
Answer #1

In Current year income assessee should adjust loss.

Building Loss= 90000(as per books)-70000(Recovery)=20000

Equipment Loss= 40000(as per books)-25000=15000

Parking Structure Loss= 90000-(70000+25000)=(5000)

So Net loss was $30000 which will be set off against business income of $220000

Net Income will be=220000-30000=190000